OpenStudy (anonymous):

Which property should be used next in this solution process? 3x + 2 + 3 = 7( x-1) - 4 3x + 5 = 7(x-1) - 4

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Which do you think?

OpenStudy (anonymous):

Distributive

OpenStudy (anonymous):

Are there two equations? 3x + 2 + 3 = 7( x-1) - 4 3x + 5 = 7(x-1) - 4

OpenStudy (anonymous):

They are the same formula. So the question is in regards to? 3x + 5 = 7(x-1) - 4

OpenStudy (anonymous):

is it not ditributive

OpenStudy (anonymous):

The question is not clear. Are you referring to that equation?

OpenStudy (anonymous):

yes both

OpenStudy (anonymous):

Why do you say distributive?

OpenStudy (anonymous):

because u have to distribute 7 to x n to - 1

OpenStudy (anonymous):

n stands for and

OpenStudy (anonymous):

Ok so if evaluate the characteristics of each property then why do you question it. You are right it is distributive.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Commutative property: When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4 Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Multiplicative Identity Property: The product of any number and one is that number. For example 5 * 1 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3